OpenStudy (anonymous):

what is (2 sq. root of 6z+1) ^2 ? My brain is dead! I would actually write it out if I could, but I don't know how to do that on here

jhonyy9 (jhonyy9):

sqrt(6z+1)^2 =6z+1

OpenStudy (amistre64):

\[\sqrt{(a*a)}=a\] \[(\sqrt{a})(\sqrt{a})=a\]

OpenStudy (anonymous):

thank you!

jhonyy9 (jhonyy9):

my pleasure good luck bye